OpenStudy (anonymous):

A right triangle has one side of length 7 and a hypotenuse of length 10. How long is the third side of the triangle?

OpenStudy (anonymous):

do pythagorean theorem

OpenStudy (anonymous):

\[a^2+b^2=c^2\]

OpenStudy (anonymous):

a and b = legs or sides of the triangle and c= hypotenuse

OpenStudy (anonymous):

\[\sqrt{51}\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

solve the \[\sqrt{51}\] u have the ur answer